Suppose a firm finds that the marginal product of capital is 60 and the marginal product of labor is 20. If the price of capital is $6 and the price of labor is $2.50, describe how the firm should adjust its mix of capital and labor? What will be the result?

1
Expert's answer
2021-11-28T17:54:21-0500

Given that;

marginal product of capital(MPK)=60

marginal product of labor (MPL)=20

wage rate(w)="\$2.50"

price of capital(r)="\$6"


for a firm employing labor and capital as input then best combination occurs at;


"\frac{MPL}{W}=\frac{MPK}{r}"


so

"\frac{MPL}{W}=\frac{20}{2.50}\\\\\\"


"\frac{MPL}{W}=8"


and


"\frac{MPK}{r}=\frac{60}{6
}"


"\frac{MPK}{r}=10"


so we can see that


"\frac{MPL}{W} <\frac{MPK}{r}\\"


"8<10"


To attain the best input mix , the firm should hire more labor and hire less capital so that marginal product of labor increases and marginal product of capital reduces until the equilibrium

is attained.